We wish to test the hypothesis that ?≤1000 at the 5% level using a sample of 25 observations. what is the critical value for this test?
Answer #1

n = 25

df = n-1 = 24

It is a right tailed test.

Right tailed critical value, t-crit = ABS(T.INV(0.05, 24)) = 1.711